OpenStudy (anonymous):

It takes 15 minutes to walk to the mall from your house at an average velocity of 1.1 m/s to the east. What is the displacement from your house to the mall?

OpenStudy (anonymous):

are you from pc?

OpenStudy (anonymous):

not to be weird but there is a football team named that where i live

OpenStudy (anonymous):

Ok.. let's see.. We know that we walked for 15 minutes, avg speed 1.1m/s so.. length=1.1m/s⋅15minutes=1.1m/s⋅(60⋅15)s=1.1m/s⋅90s=99m